# Julia in HEP

### v2

@@ 2.2 Julia in Practice
For example, the listing in Figure 1 shows some toy code to generate a grayscale image of the famous MandelbrÃt set.
{minted}
[linenos]julia using Images
function mandel(z; maxiter=80) c = z for n = 1:maxiter if abs2(z) > 4 return n-1 end z = z^2 + c end maxiter end
set = [ mandel(complex(r,i)) for i=-1.:.01:1., r=-2.0:.01:0.5 ] img = Gray.(set ./ 80)
@@ 2.3.1 Type System
{minted}
julia Float64 <: AbstractFloat <: Real <: Number <: Any
@@ 2.3.2 Multiple dispatch
{minted}
[fontsize=]julia using Measurements, Plots using DifferentialEquations
g = 9.79 Â 0.02; L = 1.00 Â 0.01; u0 = [0.0 Â 0.0, Ï  60 Â 0.01] tspan = (0.0 Â 0, 1.0 Â 0) function p(du,u,p,t) du[1]=u[2] du[2]=-(g/L)*u[1] end prob = ODEProblem(p, u0, tspan) sol = solve(prob, Tsit5()); plot(sol.t, getindex.(sol.u,2))
@@ 3.1 GPU Programming
{minted}
julia using CUDA
a = CuArray([1,2,3,4]) a * 2
@@ 3.2 Julia HPC codes
More recently, Julia has been advancing multi-GPU and multi-CPU applications in large-scale simulations. Notable examples include FastIce.jl for high-resolution flow simulations and various geocomputing applications leveraging HPC-related packages like ParallelStencils.jl and ImplicitGlobalGrid.jl. These packages enable high-performance stencil computations and efficient distributed memory parallelism, reinforcing Juliaâs role as a modern HPC language.
@@ 4.3 Event Generators
With the open-source framework QuantumElectrodynamics.jl qedjl-github , the first steps in exploring how Julia can address these challenges are taken. Specifically, the framework facilitates the numerical calculation of scattering amplitudes and the implementation of Monte Carlo event generation within the domain of perturbative and strong-field quantum electrodynamics (cf. Fedotov:2022ely ). The overall structure of the framework is unified through interfaces defined in QEDbase.jl, which provides standardised representations for fundamental mathematical objects: four-momenta, bi-spinors, and phase space points. Additionally, these interfaces extend to the configuration entities: scattering processes, computational models, and phase space layouts. Further interfaces are provided for computable quantities, such as differential cross-sections, and various samplers for Monte Carlo event generation. Here, Juliaâs multiple dispatch mechanism proves particularly powerful, allowing different implementations to be seamlessly integrated without explicitly specifying types within the interface definitions. Moreover, the domain-specific language facilitated by these interfaces – covering elements such as processes, models, and phase space layouts – combined with multiple dispatch enables the straightforward incorporation of analytical formulas whenever they are available. This typically reduces to implementing a single method for a specific function signature.
@@ 4.7 End-to-end Computing
The LEGEND LEGEND:2017AIPC experiment demonstrates how Julia can be used as a basis for end-to-end analysis in a larger physics experiment with multiple subsystems. LEGEND uses Julia as its official secondary software stack, both to verify the results of the primary software stack (written in Python) and as a test-bed for future software technologies. The whole data analysis chain, encompassing raw waveform data signal processing, ML-based data quality cuts, data calibration, event building and high-level statistical analysis is implemented completely in Julia. LEGEND uses the Bayesian Analysis Toolkit in Julia (BAT.jlSchulz:2021BAT as its primary Bayesian framework for both background decomposition and final physics analysis, and the Julia package SolidStateDetectors.jl Abt:2021SSD for detector simulation and detector design. With the exception of a custom Geant4-based software, the LEGEND collaboration is now able to perform any simulation and analysis task in Julia.

### v3

@@ 2.2 Julia in Practice
For example, the listing in Figure 1 shows some toy code to generate a grayscale image of the famous Mandelbrt set.
using Images
[…]
c = z
for n = 1:maxiter
if abs2(z) > 4
return n-1
end
z = z^2 + c
end
maxiter
end
[…]
img = Gray.(set ./ 80)
@@ 2.3.1 Type System
Float64 <: AbstractFloat <: Real <: Number <: Any
@@ 2.3.2 Multiple dispatch
using Measurements, Plots
using DifferentialEquations
[…]
L = 1.00 Âś 0.01;
u0 = [0.0 Âś 0.0, ÏĂ \ 60 Âś 0.01]
tspan = (0.0 Âś 0, 1.0 Âś 0)
function p(du,u,p,t)
du[1]=u[2]
du[2]=-(g/L)*u[1]
end
prob = ODEProblem(p, u0, tspan)
sol = solve(prob, Tsit5());
plot(sol.t, getindex.(sol.u,2))
@@ 3.1 GPU Programming
using CUDA
[…]
a * 2
@@ 3.2 Julia HPC codes
More recently, Julia has been advancing multi-GPU and multi-CPU applications in large-scale simulations. Notable examples include FastIce.jl for high-resolution flow simulations and various geocomputing applications leveraging HPC-related packages like ParallelStencils.jl and ImplicitGlobalGrid.jl. These packages enable high-performance stencil computations and efficient distributed memory parallelism, reinforcing Julias role as a modern HPC language.
@@ 4.3 Event Generators
With the open-source framework QuantumElectrodynamics.jl qedjl-github , the first steps in exploring how Julia can address these challenges are taken. Specifically, the framework facilitates the numerical calculation of scattering amplitudes and the implementation of Monte Carlo event generation within the domain of perturbative and strong-field quantum electrodynamics (cf. Fedotov:2022ely ). The overall structure of the framework is unified through interfaces defined in QEDbase.jl, which provides standardised representations for fundamental mathematical objects: four-momenta, bi-spinors, and phase space points. Additionally, these interfaces extend to the configuration entities: scattering processes, computational models, and phase space layouts. Further interfaces are provided for computable quantities, such as differential cross-sections, and various samplers for Monte Carlo event generation. Here, Julias multiple dispatch mechanism proves particularly powerful, allowing different implementations to be seamlessly integrated without explicitly specifying types within the interface definitions. Moreover, the domain-specific language facilitated by these interfaces – covering elements such as processes, models, and phase space layouts – combined with multiple dispatch enables the straightforward incorporation of analytical formulas whenever they are available. This typically reduces to implementing a single method for a specific function signature.
@@ 4.7 End-to-end Computing
The LEGEND LEGEND:2017AIPC experiment demonstrates how Julia can be used as a basis for end-to-end analysis in a larger physics experiment with multiple subsystems. LEGEND uses Julia as its official secondary software stack, both to verify the results of the primary software stack (written in Python) and as a test-bed for future software technologies. The whole data analysis chain, encompassing raw waveform data signal processing, ML-based data quality cuts, data calibration, event building and high-level statistical analysis is implemented completely in Julia. LEGEND uses the Bayesian Analysis Toolkit in Julia (BAT.jl) Schulz:2021BAT as its primary Bayesian framework for both background decomposition and final physics analysis, and the Julia package SolidStateDetectors.jl Abt:2021SSD for detector simulation and detector design. With the exception of a custom Geant4-based software, the LEGEND collaboration is now able to perform any simulation and analysis task in Julia.